OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 my answer is slope = 5/3; y-intercept = (0, -6) Am I right? :)

OpenStudy (anonymous):

5x-3y=-6 -3y=-5x-6 y=5/3x+2 The slope is right. Check your y-intercept again.

OpenStudy (anonymous):

y-intercept = (0, 2) ?

OpenStudy (anonymous):

Yes. Are you familiar with slope-intercept form?

OpenStudy (anonymous):

Sort of, I'm just starting to learn it :)

OpenStudy (anonymous):

Thanks for the help! :)

OpenStudy (anonymous):

y=mx+b m is your slope and b is going to be your y-intercept.

OpenStudy (anonymous):

No problem. Take care!

OpenStudy (anonymous):

Oh ok thanks! =D